OpenStudy (compassionate):

An aceptabe slope for a stairway is 8 inches of rise for eac 12 inches of tread or run. A builder wants to build a house with a distance of 9 feet 6 inches between the first floor and second floor. How many linear feet of room will the stairway occupy on the first loor?

OpenStudy (compassionate):

@Miracrown

OpenStudy (anonymous):

It was a little hard to visualize. I think this is what they mean 

OpenStudy (anonymous):



OpenStudy (compassionate):

I suppose they want a scale factor or something along those lines

OpenStudy (anonymous):

I think the height's 9.5 and they want the length. and the triangles are similar so we can set up an equation

Miracrown (miracrown):

Hi

OpenStudy (anonymous):

corresponding sides are proportional so \[\frac{ 12 }{ x }=\frac{ 8 }{ 9.5 }\]

OpenStudy (compassionate):

So then we cross multiply and solve, right

OpenStudy (compassionate):

144 = 8x, x = 18

OpenStudy (anonymous):

12*9.5 = 114

OpenStudy (compassionate):

right. Autism speaks. Don't know how I did that. Thanks a lot man!

OpenStudy (anonymous):

np

OpenStudy (the_fizicx99):

Noice 1, Comp GeeGee